# The Recurrence spectrum for dynamical systems beyond specification

### v1

@@ 1. Introduction
The aim of this paper is to extend Feng and Wu’s result [10] to a wide class of subshifts. For some dynamical systems with specification, Lau and Shu [21] considered recurrence sets and an associated spectrum replacing cylinders by Bowen balls and Hausdorff dimension by topological entropy. However, specification is a very strong property [4]. Inspired by the works of Climenhaga and Thompson [6, 7], in this paper we introduce (W’)-specification, and show that any recurrence set of a subshift with this property has full Hausdorff dimension (Theorem 1.1). We also establish an analogous statement for a wide class of interval maps (Theorem 1.2).
@@ 1.1. Recurrence spectrum for subshifts
We will consider the recurrence sets with respect to the partition
@@
Specification roughly means that one can glue together a collection of orbit segments to form one orbit. Points in the recurrence sets with prescribed properties may be constructed inductively, by replicating initial pieces of orbit segments which are constructed in previous steps. Hence, specification plays an important role in the construction and analysis of recurrence sets.
@@
The definition of (W’)-specification in Theorem 1.1 is given in §2.1. This property is a variant of (W)-specification introduced by Climenhaga and Thompson [6, 7] for the purpose of developing an ergodic theory for dynamical systems beyond specification. As in Lemma 2.3, (W’)-specification implies (W)-specification. Many of the presently known subshifts that satisfy (W)-specification also satisfy (W’)-specification. Below is a list of three classes of subshifts that satisfy the assumption of Theorem 1.1.
@@ 1.1.1. Coding spaces of interval maps
We say $T\colon[0,1]\to[0,1]$ is a piecewise monotonic map if there exist pairwise disjoint, finitely many non-degenerate subintervals $I_{0},\ldots,I_{m-1}$ of $[0,1]$ such that $\bigcup_{j=0}^{m-1}I_{j}=[0,1]$, and $T|_{{\rm int}(I_{j})}$ is strictly monotone and continuous for every $0\leq j\leq m-1$.
@@
By following orbits of $T$ over the partition $\xi_{T}$, one can code points in $[0,1]$ into symbol sequences in a subshift in $\Sigma_{m}$. This subshift is denoted by $\Sigma_{T}$, and called the coding space of $T$ (see §3.1). We say $T$ is transitive if there is $x\in[0,1]$ such that $\{T^{n}(x)\colon n\geq 0\}$ is dense in $[0,1]$. If $T$ is transitive then $\xi_{T}$ is a genearator. If moreover $T$ has positive topological entropy, then for any $\varepsilon>0$ there is a decomposition $\mathcal{L}(\Sigma_{T})=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ such that $\mathcal{G}$ has (W’)-specification and $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<\varepsilon$ (see Proposition 3.4). The coding space of any transitive piecewise monotonic map with positive topological entropy satisfies the assumption of Theorem 1.1.
@@ 1.1.2. Coded shifts
Let $v=v_{1}\cdots v_{|v|}$ and $w$ be words from $\{0,\ldots,m-1\}$ that are not the empty word. We say $w$ is a prefix of $v$ if $w=v_{1}\cdots v_{|v|-i}$ for some $0\leq i\leq|v|-1$. We say $w$ is a suffix of $v$ if $w=v_{i}\cdots v_{|v|}$ for some $1\leq i\leq|v|$. For $1\leq n\leq|v|$, let $v|^{\rm p}_{n}$ and $v|^{\rm s}_{n}$ denote the prefix and suffix of $v$ of length $n$ respectively. Let $\Sigma$ be a subshift for which there are subsets $\mathcal{C}^{\rm p}$, $\mathcal{G}$, $\mathcal{C}^{\rm s}$ of $\mathcal{L}(\Sigma)$ such that:
@@
    there is $t\in\mathbb{N}$ such that for all $v$, $w\in\mathcal{G}$, all $1\leq k\leq|v|$ and all $1\leq|n|\leq|w|$ there is $u\in\mathcal{L}(\Sigma)$ such that $v|^{\rm s}_{k}u(w|^{\rm p}_{n})\in\mathcal{G}$ and $|u|\leq t$;
@@
We prove (1.2) in §2 by constructing for each $(a,b)$ Moran fractals of large Hausdorff dimension contained in $R_{\sigma,\xi}(a,b)$. Moran fractals are sets that are constructed by iterative procedure using a fixed number of parameters, see e.g., [30]. The Hausdorff dimension of Moran fractals can be estimated from below with the mass distribution principle [9].
The construction of our Moran fractals breaks into three steps. In the first step, for all sufficiently large integer $k$ we construct a seed set $F_{k}$ that is independent of $(a,b)$, consists of non-recurrent points and has large Hausdorff dimension. In the second step, we modify the seed sets into subsets of the recurrence set $R_{\sigma,\xi}(a,b)$. Finally we refine these subsets to obtain our Moran fractals. Although inspired by the construction of Feng and Wu for the full shift [10], this three-step construction of Moran fractals necessarily gets more involved as developed below.
In the first step we construct the seed set $F_{k}$ by representing an ergodic measure of large entropy with a collection of words and concatenating them. Due to the existence of forbidden words, it is not possible to fully concatenate a given collection of words. (W)-specification [6, 7] plays a crucial role. A similar construction was successfully undertaken in [37] for different purposes.
(W’)-specification plays a crucial role in the second step. In §2.1 we briefly explain how it is used and why (W)-specification does not suffice. To keep the reasonable length of this paper, we only give a proof of Theorem 1.1 for one-sided shifts. A proof for two-sided shifts is completely analogous.
@@ 1.2. Recurrence spectrum for piecewise expanding interval maps
We also establish a statement analogous to Theorem 1.1 on the recurrence sets for a wide class of interval maps. We say a piecewise monotonic map $T$ is of class $C^{r}$ $(r\geq 1)$ if for each $I\in\xi_{T}$ there is an open interval $J$ containing $I$ such that $f|_{I}$ can be extended to a $C^{r}$ function on $J$. We say a piecewise monotonic map $T$ is piecewise expanding if it is of class $C^{1}$ and $\inf|T^{\prime}|>1$. If $T$ is a piecewise expanding, then $\xi_{T}$ is a generator.
@@
Theorem 1.2 is not a direct translation of Theorem 1.1. Our main tool is a Markov diagram introduced by Hofbauer [12]. In §3, we show that the coding space of any map in Theorem 1.2 satisfies the assumption in Theorem 1.1. This allows us to immediately transfer the Moran fractals in the coding space constructed in §2 to the interval $[0,1]$. When applying the mass distribution principle to the transferred Moran fractals, the difficulty is that the lengths of cylinders in the coding space are not well comparable to the Euclidean lengths of the corresponding intervals in $[0,1]$. This is due to the lack of specification or Markovianness, see e.g., [23, Proposition 2.6] for the case of the beta transformations. We have managed to obtain a nice lower bound on the lengths of intervals, see Lemma 3.5.
The rest of this paper consists of two sections. Section 2 is the analysis of recurrence sets for subshifts in Theorem 1.1. Section 3 is the analysis of recurrence sets for interval maps in Theorem 1.2.
@@ 2. Dimension of recurrence sets for subshifts
This section is devoted to the analysis of recurrence sets for subshifts in Theorem 1.1. In §2.1 we introduce (W’)-specification, and investigate its relationship with (W)-specification in [6, 7]. In §2.2 we state one lemma the proof of which is deferred to Appendix A. In §2.3 we construct seed sets. In §2.4 we modify the seed sets into recurrence sets, and in §2.5 construct Moran fractals refining these subsets. In §2.6 we complete the proof of Theorem 1.1.
2.1. Specification
We say a subshift $\Sigma$ has the periodic specification property if there is an integer $t\geq 0$ such that for every integer $k\geq 2$ and all $v^{1},\ldots,v^{k}\in\mathcal{L}(\Sigma)$, there are $w^{1},\ldots,w^{k}\in\mathcal{L}(\Sigma)$ such that $v^{1}w^{1}v^{2}w^{2}\cdots v^{k}w^{k}\in\mathcal{L}(\Sigma)$ and $|w^{i}|=t$ for $i=1,\ldots,k$, and the cylinder $v^{1}w^{1}v^{2}w^{2}\cdots v^{k}w^{k}]$ contains a periodic point. If we drop the existence of a periodic point from the above definition, then we say $\Sigma$ has the specification property. If we further replace $|w^{i}|=t$ by $|w^{i}|\leq t$, we say $\mathcal{L}(\Sigma)$ has (W)-specification ([6, p.792 Second remark]).
The periodic specification property is a shift space version of the definition of specification introduced in [3]. See also [8, 15]. The specification property and the periodic specification property are equivalent [19, §2]. (W)-specification for $\mathcal{L}(\Sigma)$ is a shift space version of the definition of specification in [21].
A natural way to relax the specification property is to require that connected words be chosen from a proper subset of $\mathcal{L}(\Sigma)$.
@@ 2.1. Specification
We will construct points in designated recurrence sets by constructing longer and longer words inductively, replicating at each step prefixes of words which are constructed in the previous step. See §2.4 for details. Clearly, (W)-specification cannot be used for this construction. Below we introduce another specification that is appropriate for this inductive construction.
@@ 2.3. Construction of seed sets
For a subshift $\Sigma$, a function $\varphi\colon\Sigma\to\mathbb{R}$ and an integer $k\geq 1$, write $S_{k}\varphi$ for the sum $\sum_{i=0}^{k-1}\varphi\circ\sigma^{i}$. Let $v$, $w=w_{1}\cdots w_{n}$ be words from $\{0,\ldots,m-1\}$. We say $v$ is a subword of $w$ if there exist $i,j\in\{1,\ldots n\}$ such that $i\leq j$ and $w_{i}\cdots w_{j}=v$. The next proposition is central to the construction of seed sets.
@@
Let $\Sigma$ be a subshift that has a language decomposition $\mathcal{L}(\Sigma)=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ with $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<h_{\rm top}(\Sigma)$. Let $\mu\in\mathcal{M}(\Sigma,\sigma)$ be ergodic and satisfy $h_{\mu}(\sigma)>h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})$. Let $\varphi\colon\Sigma\to\mathbb{R}$ be a continuous function. For any sufficiently small $\varepsilon\in(0,h_{\mu}(\sigma))$ and any $M\geq 1$, there exist an integer $k\geq M$ and a subset $\mathcal{Q}_{k}$ of $\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ such that:
@@
Step 1: construction of $\mathcal{Q}_{k}$. We use the next lemma in [37] that can be proved combining Shannon-McMillan-Breiman’s theorem and Birkhoff’s ergodic theorem. Recall the notation for a prefix of a word in §1.1.3.
@@
For each sufficiently large integer $n$, let $\mathcal{P}_{n}$ be the subset of $\mathcal{L}_{n}(\Sigma)$ given by Lemma 2.7 with $\varepsilon$ replaced by $\varepsilon^{3}$. For each $w\in\mathcal{P}_{n}$ we fix a decomposition
@@
Let $\|\varphi\|$ denote the supremum norm of $\varphi$. For any $w\in\mathcal{P}_{k_{0}}$ such that $c(w)\in\mathcal{Q}_{k}$, take $y\in[w]$. We have
@@
Combining these two inequalities and using $k/k_{0}\geq 1-2\varepsilon^{2}/\underline{h}(\mu)$ in (2.1), we have
@@ 2.4. Construction of subsets of the recurrence set
Throughout this and the next subsections, let $\Sigma$ be a subshift that has a language decomposition $\mathcal{L}(\Sigma)=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ with $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<h_{\rm top}(\Sigma)$ and $\mathcal{G}$ has (W’)-specification with gap size $t$. Let $\mu\in\mathcal{M}(\Sigma,\sigma)$ be ergodic with $h_{\mu}(\sigma)>h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})$, let $\varphi\colon\Sigma\to\mathbb{R}$ be a continuous function, let $k>3t$ be a sufficiently large integer, and let $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$, $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ be given by Proposition 2.5.
Let $a,b\in[0,\infty]$ satisfy $a\leq b$. We construct a subset of the recurrence set $R_{\sigma,\xi}(a,b)$ by modifying the seed set $F_{k}$ constructed in §2.3. Let $(\ell_{p})_{p=1}^{\infty}$, $(\gamma_{p})_{p=1}^{\infty}$ be sequences of positive reals with the properties in Lemma 2.4. By virtue of (b) and (c) in Lemma 2.4, shifting the indices if necessary we may assume
@@
By (2.7) and (2.8), $(M_{p}(x))_{p=1}^{\infty}$ is well-defined. We construct a sequence $(y_{q}(x))_{q=0}^{\infty}$ of words in $\mathcal{L}(\Sigma)$ by induction. Start with $y_{0}(x)=v^{*}$. Suppose $q=0$, or $q\geq 1$ and $y_{q}(x)=v^{*}\cdots v^{q}(x)$ is a concatenation of $v^{*}$ and elements of $\mathcal{Q}_{k}$ by words of lengths not exceeding $t$. There are two cases.
@@
Eligible choices of the connecting word $w$ in Case 1 and that of $w^{p,1}$, $w^{p,2}$, $w^{p,3}$, $\lambda^{p}(x)$ in Case 2 are not unique. To avoid any ambiguity, we set a definite selection rule right from the start: introduce total orders in $\bigcup_{n=0}^{t}\mathcal{L}_{n}(\Sigma)$ and $\mathcal{Q}_{k}$, and choose from those eligible words the ones that are minimal with respect to these orders. This completes the construction of the sequence $(y_{q}(x))_{q=0}^{\infty}$. Let $y(x)$ denote the element of the singleton $\bigcap_{q=0}^{\infty}[y_{q}(x)]$. Clearly we have $y(x)\in\Sigma$. Let $\Lambda_{k}(a,b)$ denote the collection of these points:
@@
By (2.6) and (2.11), the sequence $(\theta^{p}(x))_{p=1}^{\infty}$ is strictly monotone increasing.
$\tau_{\xi_{|\theta^{p}(x)|}(y)}(y)=|y_{M_{p}(x)}(x)w^{p,1}|.$
@@
$e^{\gamma_{p}\ell_{p}}\leq\tau_{\xi_{|\theta^{p}(x)|}(y(x))}(y(x))\leq(1+\eta)e^{\gamma_{p}\ell_{p}}+k+t.$
This yields the desired equality. ∎
@@
Let $x\in F_{k}$ and write $y$ for $y(x)$. Combining (2.11), Lemma 2.4(a) and Lemma 2.8 we get
(2.12) $\lim_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{2p+1}(x)|}(y)}(y)}{|\theta^{2p+1}(x)|}=a\ \text{ and }\ \lim_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{2p}(x)|}(y)}(y)}{|\theta^{2p}(x)|}=b.$
@@
(2.13) $\limsup_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y)}(y)}{|\theta^{p}(x)|}=\limsup_{p\to\infty}\frac{\gamma_{p+1}\ell_{p+1}}{\ell_{p}}\leq b.$
@@
$s_{p}(x)=\max\{i\geq|\theta^{p}(x)|\colon(\theta^{p}(x)w^{p,2}\lambda^{p}(x))|^{\rm p}_{i}\vartriangleleft y_{M_{p}(x)}(x)\}.$
Notice that $\mathbb{N}\setminus\{1,\ldots,\theta^{1}(x)-1\}=\bigsqcup_{p\in\mathbb{N}}(N_{1}(p)\bigsqcup N_{2}(p))$. By the choice of $\lambda^{p}(x)$, the number $s_{p}(x)$ is well-defined and satisfies

### v2

@@ 1. Introduction
The aim of this paper is to extend Feng and Wu’s result [10] to a wide class of subshifts. For some dynamical systems with specification (see section 2.1 for various notions of specification), Lau and Shu [21] considered recurrence sets and an associated spectrum replacing cylinders by Bowen balls and Hausdorff dimension by topological entropy. However, the notion of specification used in [21] is strong. Inspired by the works of Climenhaga and Thompson [6, 7], in this paper we introduce (W’)-specification, and show that any recurrence set of a subshift with (W’)-specification has full Hausdorff dimension (theorem 1.1). We also establish an analogous statement for a wide class of interval maps (theorem 1.2).
@@ 1.1. Recurrence spectrum for subshifts
We will consider the recurrence sets in $\Sigma$ with respect to the natural partition
@@
We say a subshift $\Sigma$ has the specification property if there is an integer $t\geq 0$ such that for all $u,v\in\mathcal{L}(\Sigma)$ there is $w\in\mathcal{L}(\Sigma)$ such that $uwv\in\mathcal{L}(\Sigma)$ and $|w|=t$. This property roughly means that one can glue together a collection of orbit segments to form one orbit. Points in the recurrence sets with prescribed properties may be constructed inductively, by replicating initial pieces of orbit segments which are constructed in previous steps. Hence, the specification property plays an important role in the analysis of recurrence sets.
@@
The definition of (W’)-specification in theorem 1.1 is given in section 2.1. This property is a variant of (W)-specification introduced by Climenhaga and Thompson [6, 7] for the purpose of developing an ergodic theory for dynamical systems beyond specification. As in lemma 2.3, (W’)-specification implies (W)-specification. Many of the presently known subshifts that satisfy (W)-specification also satisfy (W’)-specification. Below is a list of three classes of subshifts that satisfy the assumption of theorem 1.1.
@@ 1.1.1. Coding spaces of interval maps
We say $T\colon[0,1]\to[0,1]$ is a piecewise monotonic map if there exist pairwise disjoint, finitely many non-degenerate subintervals $I_{0},\ldots,I_{m-1}$ of $[0,1]$ such that $\bigcup_{j=0}^{m-1}I_{j}=[0,1]$, and the restriction $T|_{{\rm int}(I_{j})}$ is strictly monotone and continuous for every $0\leq j\leq m-1$.
@@
By following orbits of $T$ over the partition $\xi_{T}$, one can code points in $[0,1]$ into symbol sequences in a subshift in $\Sigma_{m}$. This subshift is denoted by $\Sigma_{T}$, and called the coding space of $T$ (see section 3.1). We say $T$ is transitive if there is $x\in[0,1]$ such that $\{T^{n}(x)\colon n\geq 0\}$ is dense in $[0,1]$. If $T$ is transitive then $\xi_{T}$ is a genearator. If moreover $T$ has positive topological entropy, then for any $\varepsilon>0$ there is a decomposition $\mathcal{L}(\Sigma_{T})=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ such that $\mathcal{G}$ has (W’)-specification and $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<\varepsilon$ (see proposition 3.4). The coding space of any transitive piecewise monotonic map with positive topological entropy satisfies the assumption of Theorem 1.1.
@@ 1.1.2. Coded shifts
Let $v=v_{1}\cdots v_{|v|}$ and $w$ be words from $\{0,\ldots,m-1\}$ that are not the empty word. We say $w$ is a prefix of $v$ if $w=v_{1}\cdots v_{|v|-i}$ for some $0\leq i\leq|v|-1$. We say $w$ is a suffix of $v$ if $w=v_{i}\cdots v_{|v|}$ for some $1\leq i\leq|v|$. For each $n\in\{1,\ldots,|v|\}$, let $v|^{\rm p}_{n}$ and $v|^{\rm s}_{n}$ denote the prefix and suffix of $v$ of length $n$ respectively. Let $\Sigma$ be a subshift for which there are subsets $\mathcal{C}^{\rm p}$, $\mathcal{G}$, $\mathcal{C}^{\rm s}$ of $\mathcal{L}(\Sigma)$ such that:
@@
    there is $t\in\mathbb{N}$ such that for all $v$, $w\in\mathcal{G}\setminus\{\emptyset\}$ there is $u\in\mathcal{L}(\Sigma)$ with $|u|\leq t$ such that $v|^{\rm s}_{k}u(w|^{\rm p}_{n})\in\mathcal{G}$ whenever $k,n\in\mathbb{N}$ satisfy $1\leq k\leq|v|$, $1\leq n\leq|w|$ and $v|^{\rm s}_{k}\in\mathcal{G}$, $w|^{\rm p}_{n}\in\mathcal{G}$;
@@
We prove (1.2) in section 2 by constructing for each $(a,b)$ Moran fractals of large Hausdorff dimension contained in $R_{\sigma,\xi}(a,b)$. Moran fractals are sets that are constructed by iterative procedure using a countable number of parameters, see e.g., [30]. The Hausdorff dimension of Moran fractals can be estimated from below with the mass distribution principle [9].
The construction of our Moran fractals breaks into three steps. In the first step (section 2.3), for all sufficiently large integer $k$ we construct a seed set $F_{k}$ that is independent of $(a,b)$, consists of non-recurrent points and has large Hausdorff dimension. In the second step (section 2.4), we modify the seed sets into subsets of the recurrence set $R_{\sigma,\xi}(a,b)$. In the final step (section 2.5) we refine these subsets to obtain our Moran fractals. Although inspired by the construction of Feng and Wu for the full shift [10], our construction of Moran fractals necessarily gets more involved as developed below.
In the first step we construct the seed set $F_{k}$ by representing an ergodic measure of large entropy with a collection of words and concatenating them. Due to the existence of forbidden words, it is not possible to fully concatenate a given collection of words, and (W)-specification [6, 7] plays a crucial role. A similar construction was successfully undertaken in [37] for different purposes.
(W’)-specification plays a crucial role in the second step. In section 2.1 we briefly explain how it is used and why (W)-specification does not suffice. To keep the reasonable length of this paper, we only give a proof of theorem 1.1 for one-sided shifts. A proof for two-sided shifts is completely analogous.
@@ 1.2. Recurrence spectrum for piecewise expanding interval maps
We also establish a statement analogous to theorem 1.1 on the recurrence sets for a wide class of interval maps. We say a piecewise monotonic map $T$ is of class $C^{r}$ $(r\geq 1)$ if for each $I\in\xi_{T}$ there is an open interval $J$ containing $I$ such that $T|_{I}$ can be extended to a $C^{r}$ function on $J$. We say a piecewise monotonic map $T$ is piecewise expanding if it is of class $C^{1}$ and $\inf|T^{\prime}|>1$. If $T$ is a piecewise expanding, then $\xi_{T}$ is a generator.
@@
Theorem 1.2 is not a direct translation of theorem 1.1. Our main tool is a Markov diagram introduced by Hofbauer [12]. In section 3, we show that the coding space of any map in theorem 1.2 satisfies the assumption in theorem 1.1. This allows us to immediately transfer the Moran fractals in the coding space constructed in section 2 to the interval $[0,1]$. When applying the mass distribution principle to the transferred Moran fractals, the difficulty is that the lengths of cylinders in the coding space are not well comparable to the Euclidean lengths of the corresponding intervals in $[0,1]$. This is due to the lack of specification or Markovianness, see e.g., [23, proposition 2.6] for the case of the beta transformations. We have managed to obtain a nice lower bound on the lengths of intervals, see lemma 3.5.
The rest of this paper consists of two sections. Section 2 is devoted to the analysis of recurrence sets for subshifts in theorem 1.1. Section 3 is devoted to the analysis of recurrence sets for interval maps in theorem 1.2.
@@ 2. Dimension of recurrence sets for subshifts
This section is devoted to the analysis of recurrence sets for subshifts in theorem 1.1. In section 2.1 we introduce (W’)-specification, and investigate its relationship with (W)-specification in [6, 7]. In section 2.2 we state one lemma the proof of which is deferred to appendix A. In section 2.3 we construct seed sets. In section 2.4 we modify the seed sets into recurrence sets, and in section 2.5 construct Moran fractals refining these subsets. In section 2.6 we complete the proof of theorem 1.1.
2.1. Various notions of specification
We say a subshift $\Sigma$ has the periodic specification property if there is an integer $t\geq 0$ such that for every integer $k\geq 2$ and all $v^{1},\ldots,v^{k}\in\mathcal{L}(\Sigma)$, there are $w^{1},\ldots,w^{k}\in\mathcal{L}(\Sigma)$ such that $v^{1}w^{1}v^{2}w^{2}\cdots v^{k}w^{k}\in\mathcal{L}(\Sigma)$ and $|w^{i}|=t$ for $i=1,\ldots,k$, and the cylinder $v^{1}w^{1}v^{2}w^{2}\cdots v^{k}w^{k}]$ contains a periodic point of period $|v^{1}w^{1}v^{2}w^{2}\cdots v^{k}w^{k}|$. If we further replace $|w^{i}|=t$ by $|w^{i}|\leq t$, we say $\mathcal{L}(\Sigma)$ has (W)-specification ([6, p.792 second remark]).
The periodic specification property is a shift space version of the definition of specification introduced in [3]. See also [8, 15]. The specification property and the periodic specification property are equivalent [19, section 2]. The (W)-specification for $\mathcal{L}(\Sigma)$ is a shift space version of the definition of specification in [21].
A natural way to relax the specification property is to require that connected words be chosen from a proper subset of $\mathcal{L}(\Sigma)$ as follows.
@@ 2.1. Various notions of specification
We will construct points in designated recurrence sets by constructing longer and longer words inductively, replicating at each step prefixes of words which are constructed in the previous step. See section 2.4 for details. Clearly, (W)-specification cannot be used for this construction. Below we introduce another specification that is appropriate for this inductive construction.
@@ 2.3. Construction of seed sets
Let $\Sigma$ be a subshift. Let $\mathcal{M}(\Sigma,\sigma)$ denote the set of $\sigma$-invariant Borel probability measures on $\Sigma$. For a function $\varphi\colon\Sigma\to\mathbb{R}$ and an integer $k\geq 1$, write $S_{k}\varphi$ for the sum $\sum_{i=0}^{k-1}\varphi\circ\sigma^{i}$. Let $v$, $w=w_{1}\cdots w_{n}$ be words from $\{0,\ldots,m-1\}$. We say $v$ is a subword of $w$ if there exist $i,j\in\{1,\ldots n\}$ such that $i\leq j$ and $w_{i}\cdots w_{j}=v$. The next proposition is central to the construction of seed sets.
@@
Let $\Sigma$ be a subshift that has a language decomposition $\mathcal{L}(\Sigma)=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ with $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<h_{\rm top}(\Sigma)$. Let $\mu\in\mathcal{M}(\Sigma,\sigma)$ be ergodic and satisfy $h_{\mu}(\sigma)>h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})$. Let $\varphi\colon\Sigma\to\mathbb{R}$ be a continuous function. For any sufficiently small $\varepsilon\in(0,h_{\mu}(\sigma))$ and any $M\geq 1$, there exist an integer $k\geq M$, a subset $\mathcal{Q}_{k}$ of $\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ such that:
@@
Step 1: construction of $\mathcal{Q}_{k}$. We use the next lemma in [37] that can be proved combining Shannon-McMillan-Breiman’s theorem and Birkhoff’s ergodic theorem. Recall the notation for a prefix and suffix of a word in section 1.1.3.
@@
For each sufficiently large $n\in\mathbb{N}$, let $\mathcal{P}_{n}$ denote the subset of $\mathcal{L}_{n}(\Sigma)$ given by lemma 2.7 with $\varepsilon$ replaced by $\varepsilon^{3}$. For each $w\in\mathcal{P}_{n}$ we fix a decomposition
@@
Let $\|\varphi\|=\sup_{x\in\Sigma}|\varphi(x)|$. For any $w\in\mathcal{P}_{k_{0}}$ such that $c(w)\in\mathcal{Q}_{k}$, take $y\in[w]$. Using the equality and the first inequality in (2.1), we have
@@
Combining these two inequalities and using the last inequality in (2.1), we have
@@ 2.4. Construction of subsets of the recurrence set
Throughout this and the next subsections, let $\Sigma$ be a subshift that has a language decomposition $\mathcal{L}(\Sigma)=\mathcal{C}^{\rm p}\mathcal{G}\mathcal{C}^{\rm s}$ with $h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})<h_{\rm top}(\Sigma)$ and $\mathcal{G}$ has (W’)-specification with gap size $t$. Let $\mu\in\mathcal{M}(\Sigma,\sigma)$ be ergodic with $h_{\mu}(\sigma)>h(\mathcal{C}^{\rm p}\cup\mathcal{C}^{\rm s})$, let $\varphi\colon\Sigma\to\mathbb{R}$ be a continuous function, and let $k>\max\{3t,1\}$ be a sufficiently large integer for which there exist $\mathcal{Q}_{k}\subset\mathcal{G}_{k}$ and $v^{*}\in\mathcal{G}_{k}\setminus\mathcal{Q}_{k}$ as in proposition 2.5.
Let $a,b\in[0,\infty]$ satisfy $a\leq b$. We construct a subset of the recurrence set $R_{\sigma,\xi}(a,b)$ by modifying the seed set $F_{k}$ constructed in section 2.3. Let $(\ell_{p})_{p=1}^{\infty}$, $(\gamma_{p})_{p=1}^{\infty}$ be sequences of positive reals with the properties in lemma 2.4. By virtue of (b) and (c) in lemma 2.4, shifting the indices if necessary we may assume
@@
By (2.7) and (2.8), $(M_{p}(x))_{p=1}^{\infty}$ is well-defined. We construct a sequence $(y_{q}(x))_{q=0}^{\infty}$ of words in $\mathcal{L}(\Sigma)$ by induction. Start with $y_{0}(x)=v^{*}$. Suppose $q=0$, or else $q\geq 1$ and $y_{q}(x)=v^{*}\cdots v^{q}(x)$ is a concatenation of $v^{*}$ and elements of $\mathcal{Q}_{k}$ by words of lengths not exceeding $t$. There are two cases.
@@
Eligible choices of the connecting word $w$ in case 1 and that of $w^{p,1}$, $w^{p,2}$, $w^{p,3}$, $\lambda^{p}(x)$ in case 2 are not unique. To avoid any ambiguity, we set a definite selection rule right from the start: introduce total orders in $\bigcup_{n=0}^{t}\mathcal{L}_{n}(\Sigma)$ and $\mathcal{Q}_{k}$, and choose from those eligible words the ones that are minimal with respect to these orders. This completes the construction of the sequence $(y_{q}(x))_{q=0}^{\infty}$. Let $y(x)$ denote the element of the singleton $\bigcap_{q=0}^{\infty}[y_{q}(x)]$. Clearly we have $y(x)\in\Sigma$. Let $\Lambda_{k}(a,b)$ denote the collection of these points, namely
@@
By (2.6) and (2.11), the sequence $(\theta^{p}(x))_{p=1}^{\infty}$ is strictly monotone increasing and
$\tau_{\xi_{|\theta^{p}(x)|}(y(x))}(y(x))=|y_{M_{p}(x)}(x)w^{p,1}|.$
@@
$e^{\gamma_{p}\ell_{p}}\leq\tau_{\xi_{|\theta^{p}(x)|}(y(x))}(y(x))\leq(1+\eta)e^{\gamma_{p}\ell_{p}}+k+t,$
which implies the desired equality. ∎
@@
Let $x\in F_{k}$. Combining (2.11), lemma 2.4(a) and lemma 2.8 we get
(2.12) $\lim_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{2p+1}(x)|}(y(x))}(y(x))}{|\theta^{2p+1}(x)|}=a\ \text{ and }\ \lim_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{2p}(x)|}(y(x))}(y(x))}{|\theta^{2p}(x)|}=b.$
@@
(2.13) $\limsup_{p\to\infty}\frac{\log\tau_{\xi_{|\theta^{p+1}(x)|}(y(x))}(y(x))}{|\theta^{p}(x)|}=\limsup_{p\to\infty}\frac{\gamma_{p+1}\ell_{p+1}}{\ell_{p}}\leq b.$
@@
$s_{p}(x)=\max\{i\geq|\theta^{p}(x)|\colon(\theta^{p}(x)w^{p,2}\lambda^{p}(x))|^{\rm p}_{i}\text{ is a prefix of }y_{M_{p}(x)}(x)\}.$
By the choice of $\lambda^{p}(x)$, the number $s_{p}(x)$ is well-defined and satisfies
@@
By the inequality in (2.10) and $k>1$, $\theta^{p}(x)\geq 2$ holds. Notice that
$\mathbb{N}\setminus\{1,\ldots,|\theta^{1}(x)|-1\}=\bigsqcup_{p\in\mathbb{N}}(N_{1}(p)\bigsqcup N_{2}(p)).$